OpenStudy (anonymous):

Please can anyone help me on the steps as to how to solve this problem: (a-2)^2=8

OpenStudy (perl):

did you try square root?

OpenStudy (anonymous):

I think the steps are as follows

OpenStudy (anonymous):

a-2 = 2 at the square root of 2 a-2 = -2 at the square root of 2

OpenStudy (anonymous):

a= 2+2 at the square root of 2 or a-2 = -2 at the square root of 2 final answer is this correct?

OpenStudy (perl):